Category Archives: Status

MFF Equipment Status

Instrument Status Status Notes
6″ Sputter deposition  Ready Online

Issues – Water flow meter is not indicating flow

20190511 – Installed a barrier to block water from getting to the wall instead of the drain

20180102 – Replaced the outlet water connectors as one had cracked and leaked water on the floor

20180406 – Replaced the pump oil and cleaned both the top and bottom seals on the chamber

20181014 – Replaced the O2 MFC as after 20 mins the flow would go to zero

8″ Sputter deposition Ready Online

Issues – Need to fix the hoist or take weight off the system

20190511 – Installed a barrier to block water from getting to the wall instead of the drain

Denton e-beam deposition  Unavailable Offline

Issues – Needs to be bead blasted

20190618 – Sent back the cryo compressor as it would not turn on

20190511 – Installed a barrier to block water from getting to the wall instead of the drain

20190418 – Cryo compressor rebuilt and reinstalled – system is functional

20181205 – Replaced the oil casing gasket as the rough pump was leaking oil

20180102 – Cryo Regenerated

Fredrick e-beam deposition  Useable Online

Issues – Needs to be bead blasted, new multiple sample holder installed, and redo the water interlock and possibly all the interlocks

20190618 – Replaced the white sensing wire (no sure actually what it was) on the HV controller as it had melted in the sensing coil

20190511 – Installed a barrier to block water from getting to the wall instead of the drain

20171024 – Water lines replaced after one cracked, New Al crucible

20190219 – Topped up the Helium to 250 psi (when off) as the operation pressure was 80 psi, should be 105 psi

Parylene Coater  Ready Online

20190511 – Installed a barrier to block water from getting to the wall instead of the drain

TRION  Ready Online

Issues – Need to install an automatic water shut off

20190529 – Hole in the water line before the plastic to metal tubing connector in the RF system. Replaced all the water lines, flushed the RF loop as the 50% DIW / 50% Ethylene Glycol mixture was not flowing in the RF cooling loop. Opened both RF supplies and checked for coolant. Flushed the system twice and emptied the debris filter in the pump head.

20190517 – Oil case seal replaced to resolve an oil leak, oil filtered, motor in the chiller replaced with the motor from the original chiller as the shaft was unstable

20171127 – Replaced the oil after the pump run out, leak in a seal still

20180622 – Replaced the lid o-ring (#383) there was a leak rate of 1Torr/min

20190108 – Replaced the rough pump with the one from the MARCH – rough pump time of ~60 secs

20190108 – Oil filter was jammed with dried oil contamination – rebuilt with old pump head (need to bleed to get the filter to draw in oil)

20190219 – CAIBE rough pump is now used for the TRION until the 2021 C2 pump is rebuilt (FY20)

CAIBE  Unavailable Offline

20180420 – Installed new stage, fixed the water line that melted, removed the load lock turbo from the cooling as it was blocked and not needed

20180601 – Installed a new microwave source

20181023 – Replaced the shaft seal on the rough pump

20190219 – Rough pump is being used by the TRION

March  Ready Ready

Issues – need to modify the side to allow for the RF cable to penetrate the side

20180127 – Connected water lines to the RF base

20190108 – RF supply is reading a power when it is on and set to zero, rough pump is on the TRION, RF supply is also not tuning and has major RF swings.

20190219 – Put a new copper ground from the vacuum line to the water line in the chase, without the MARCH RF goes crazy

20190315 – Replacement RF supply received and in working order. C1 is not turning properly as the bearing is bad.

20190320 – Rebuilt C1 and moved the RF input to go directly to the variable capacitors. System is tuning properly from 0W to 300W.

Ellipsometer  Ready Ready

Issues – None

20171024 – Light source realigned to give a lower gain of 2

FTIR  Ready Ready

Issues – Calibration

Curvature System  Ready Ready

Issues – Calibration

Oxidation Furnace  Ready Ready

Fall 2017 – Installed a water bubbler, but the lines are not sending the steam to the furnace correctly.

20180131 – Installed a new TC for zone 2, testing in the lab class

20180604 – Installed a new controller and the system seems stable

Diffusion Furnace  Ready Ready

Issues – None

20190201 – Installed a quartz liner for MDE to anneal their SU-8 films

2″ Furnace  Ready Ready

Issues – None

RTP  Useable Chiller is leaking water if run at too higher pressure

Issues – Need to make sure we have the proper cooling, system was run with glass samples and melted to the silicon TC and damaged it (2019 May)

Dicing Saw  Ready Ready

Fall 2017 – Moved the water supply off the tape to the building supply, replaced a leaking water pressure switch

20180413 – Replaced the pressure flow switch but it is not fast enough or not sending the right signal that there is water in time.

20180601 – Replaced several of the boards after boards 6 and 7 were swapped as recommended by the company and this lead to several issues with burnt out leads and blown fuses

20180814 – Installed a 12VDC power supply that will need to be manually turned on to turn on the water, fix for the water pressure issues. The optics head was realigned as it became lose and users had been bumping it.

20180901 – Installed a Cole-Parmer Venturi Vacuum Pump, 3.2 cfm, 28.0″Hg

EV620 Mask Aligner  Ready Ready – Turning off the lamp to preserve the life time

Issues – Logging sometimes fails as the optics stage gets stuck

20180814 – Fixed three air leaks, two on the right side of the stage as the tubes had cracked

Wet bench – Chemical processing  Ready Ready

Issues – Air leaks

Vacuum oven  Ready Ready

Issues – None

20180131 – Installed a vacuum pump

Hotplates  Ready Ready

Issues – None

Spinner  Ready Ready

Issues – None

Wire bonder  Ready Ready

Issues – None

Cleanroom Optical Microscopy  Ready Ready

Issues – Calibrate the CCD capture measurements, can use the SYP mask

20190219 – 100x objective repaired

Probe station  Ready Ready

Issues – probes need better bases

20180913 x-y belt replaced with a No slip positive drive belt RBC PIC DESIGN PART#F32BS18-112

Parameter Analyzer  Unavailable Offline

Issues – Communication errors with the SMU, system needs to be rebuilt $30K

Capacitance Voltage Measurement  Ready Ready

Issues – Need to automate

Four point conductivity  Ready Ready

Issues – Need to automate

Atomic Layer Deposition Ready Ready

20190401 – New TDMAH installed and Hf deposition tuned to give a 4″ uniform deposition, growth rate is too fast per cycle, need to calibrate for the ALD region

20181205 – Milled a trench in the sleeve, put in the old ceramic pins, put in metal pins to stop the sleeve from moving

20181015 – Redo the N2 to be SS instead of plastic

201805023 – set up the process to represent an ALD valve that has constant Ar carrier flow

20180601 – Installed Ferrocene iron ALD source

20180814 – changed the precursor setup – Ferrocene is not producing a vapor pressure?

20190102 – Replaced the manual shut off valve on the ferrocene with an all metal valve – finally do not see F – not sure if it was the valve. Disconnected all the lines to the shower head that is not used.

20190219 – Installed the TDMAH to test uniformity and optimize the system

Filmertics 3D Profiler Ready Ready

20171127 – 100x received and installed

20180814 – Had to realign the fringes with the focus point.

Big blue wet benches Useable Ready

20181015 – Connected the 3-phase supply to the left side – Alarm 1 is coming on so it is not running yet.

Connected the aspirator, ran a new waste line to the current acid neutralization system. Need to install an automatic water shut off valve to the house water as the shut off requires a 10′ ladder to get to.
1 – I put in a new controller for the right bench and it allows for the bench to be used. The old controller might be ok, but I did not realize you had to power cycle the bench with the main breaker if the EPO is pressed. There is still an alarm going and I am not sure yet what is causing it.
2 – I started checking the controls to see if everything works – did not get very far, but there is running DIW and N2
3 – The aspirators are not working and this is very much likely because Bill has decided to run these to the new acid waste storage tank instead of the current one. To finish the new tank will take significant time and funds to get operational and with the current usage I do not see the point. I might be able to tee into the current tank, however I am not sure what has been done to turn off the aspirators.
4 – The left bath seems to have no power, but the breaks are not labelled to know where the power is going – luckily for me there is only one break currently off – however it did not turn on the left controller for the bench.
5 – The left bench is not connected at all for waste and this will mean breaking the lines on the 3rd floor to install tees and then running tubes up to the 4th floor to be connected.
EasyNano CVD CNT Ready Ready

Installed in January – April 2019 and fully operational – installation included HVAC, flammable gas cabinet (H2, C2H4), gas lines for Ar, N2, H2, C2H4, 3-phase power